OpenStudy (anonymous):

1/x - 1/(x+5) = ???/x(x+5)

OpenStudy (anonymous):

I think it's 5, but I need reassurance please!

OpenStudy (anonymous):

it would be x-x+5, which is 5, correct? For the numerator?

OpenStudy (anonymous):

yes you are correct.

OpenStudy (anonymous):

5 /(x(x+5))

OpenStudy (anonymous):

Awesome! I thought it was right, but I'm getting tired and there is MORE to do! THANKS!

OpenStudy (anonymous):

\[\frac{1}{x} - \frac{1}{x+5} = \frac{1(x+5)}{x(x+5)} - \frac{1(x)}{x(x+5)} = \frac{x+5-x}{x(x+5)}=\frac{5}{x(x+5)}\]